please help me solve this equation (4y-9=15)
Answer by funmath(2933)

The goal is to get y all by itself on one side of the equal sign. We do this by adding and subtracting and multiplying and dividing the numbers to accomplish that. Remember, whatever you do to one side of the eqaution you have to do to the other.
4y-9=15
4y-9+9=15+9
4y+0=24
4y=24
4y/4=24/4
y=6
----------------------
Check your answer by substituting y=6 into the equation and see if both sides actually are equal when y=6.
4(6)-9=15
24-9=15
15=15
Yeah, we're right!!!!!
Happy calculating!